# Dipper: Diversity in Prompts for Producing Large Language Model Ensembles in Reasoning tasks

### v1

@@ Abstract
Large Language Models still encounter substantial challenges in reasoning tasks, especially for smaller models, which many users may be restricted to due to resource constraints (e.g. GPU memory restrictions). Inference-time methods to boost LLM performance, such as prompting methods to invoke certain reasoning pathways in responses, have been shown effective in past works, though they largely rely on sequential queries. The ensemble method, which consists of multiple constituent models running in parallel, is a promising approach to achieving better inference-time performance, especially given recent developments that enabled significant speed-ups in LLM batch inference. In this work, we propose a novel, training-free LLM ensemble framework where a single LLM model is fed an optimized, diverse set of prompts in parallel, effectively producing an ensemble at inference time to achieve performance improvement in reasoning tasks. We empirically demonstrate that our method leads to significant gains on math reasoning tasks, e.g., on MATH, where our ensemble consisting of a few small models (e.g., three Qwen2-MATH-1.5B-it models) can outperform a larger model (e.g., Qwen2-MATH-7B-it).
@@ 1 Introduction
While Large Language Models (LLMs) have demonstrated impressive capabilities in addressing a variety of tasks, they still encounter substantial challenges in reasoning tasks such as multi-step logical inference or problem-solving [1]. This is especially so for smaller models, which many users may be restricted to due to resource constraints (e.g. GPU memory restrictions), posing limitations on their utility in practice. Inference-time methods to boost LLM performance, especially for smaller models, hold promise in tackling these challenges [2]. However, many of these methods, such as Chain-of-Thought (CoT), Reflexion, and other techniques [3, 4, 5, 6], have focused on sequential queries to an LLM to improve performance.
@@
In contrast, ensemble methods, which involve the use of multiple constituent models in parallel, have been shown to improve models’ performance and robustness in classical machine-learning settings [7] and are promising approaches to achieve better inference-time performance, although less well-studied in the LLM setting. The prospects of applying such methods to LLMs are increasingly attractive, given recent developments that have enabled significant speed-ups in parallel, LLM batch inference. These include methods to efficiently handle key-value cache memory [8] and prompt caching to efficiently reuse common prompts for multiple queries [9, 10], enabling sub-linear (in the number of queries) costs for batch inference.
However, a key challenge in achieving high performing ensembles is how diversity can be appropriately injected among its constituents [11, 12], and this applies to LLM ensembles as well. Recent works have explored how using hetereogenous model ensembles (i.e. consisting of different models types) could lead to improved performance [13, 14], although users may often prefer to or be restricted to using only a single type of LLM model in practice, making such methods not viable in those cases. While a single LLM may be sampled with the same query multiple times and rely on the stochasticity of the LLM response generation process [15] to essentially form a self-ensemble, this approach injects limited diversity to the ensemble which may limit performance improvements.
Instead, significantly more diversity could potentially be injected into an LLM ensemble by making use of LLMs’ ability to produce diverse output for a given task with just different prompts. For example, adding simple prompt instructions on how the LLM should reason [16] has been shown to result in performance boosts. This leads to the interesting question of how we could design ensemble methods that rely on just prompt diversity to produce significant performance boost for a given LLM. Such ensemble methods could be applied during inference to improve the performance of any LLM (e.g. assessed via APIs), together with other types of inference-time methods.
In this work, we propose Dipper, a novel, training-free LLM ensemble framework where a single LLM model type is fed an optimized, diverse set of reasoning prompts in parallel, effectively producing an ensemble at inference time to improve performance in reasoning tasks. This approach is simple but surprisingly effective and efficient, and could be implemented with any black-box LLM.
2 Problem setting and related work
@@
LLMs and prompts. Consider an LLM model $M$ which for our purposes can be viewed as a black box that encodes a conditional probability distribution of text responses $y$ over any text input $q$ and additional prompt $w$, from which we can autoregressively sample responses $\hat{y}$ from, i.e.
@@
Examples of prompt $w$ could include reasoning prompts such as "Let’s think step by step" in CoT [17] that provide instructions on how the LLMs should derive answers for the query $q$.
LLM ensembles. Ensemble methods involve combining several models to produce a ensemble with better performance and lower variance. However, while commonly applied for a wide variety of machine learning models [7], ensemble methods for LLMs have remained relatively unexplored. Past works have focused on heterogeneous ensembles involving multiple types of models (e.g. different LLM API providers) [13], multi-agent LLM settings that focuses on interactions among agents [18, 19, 20], or homogeneous ensembles that rely only on stochastic sampling of model responses [15].
However, to the best of our knowledge, we are not aware of any work that focused on designing and analyzing homogeneous LLM ensembles where their diversity is injected and optimized via prompts to constituents with the same underlying LLM model. Our work’s focus on such an approach exploits LLMs’ unique capabilities of generating diverse output given only changes to its prompts, allowing for a simple but effective method to boost LLM performance using inference-time compute.
Problem formulation. Consider a task $\mathcal{T}$ that consists of instances described as tuples $t\coloneq(q_{t},c^{*}_{t})$, where $q_{t}$ can be represented as a text string query and $c^{*}_{t}$ is the corresponding ground truth solution. We have access to a single LLM model $M$ that when provided task queries and a prompt $w$, will provide a response $\hat{y}$ according to Eq. 1. This response will consist of (1) some reasoning output $\hat{r}$, and (2) the final answer $\hat{c}$ to the query, which we can denote as $\hat{y}\coloneq\{\hat{r},\hat{c}\}$. We evaluate the performance of the model with a specific prompt, denoted as $M(\cdot,w)$, on the task by computing its expected accuracy over the set of task instances $\mathcal{T}$, i.e., $F(M(\cdot,w);\mathcal{T})\coloneq\mathbf{E}_{t\sim\mathcal{T}}[\mathbb{I}\{\hat{c}_{t}=c^{*}_{t}\}]$, which in practice is computed over a representative test set.
@@
Our objective is to design an ensemble framework with an appropriate design parameter $\phi$ such that given fixed $M$, $n$ and a small labeled development set, we can efficiently maximize Eq. 2 by optimizing for $\phi$ to produce the best performing ensemble without additional training.
@@ 3 Method
Drawing inspiration from how using different prompts $w$ would result in varying response distributions in Eq. 1 given the same model $M$, our Dipper framework has the set of prompts $\{w_{i}\}_{i=1}^{n}$ fed into the ensemble of $n$ LLM instances as the key ensemble design parameter $\phi$. Dipper consists of the following three components:
@@
    Prompt Generator. First, an LLM generates a large candidate pool of prompts (denoted as $\mathcal{W}$), which can be based on some description of the task and in-context prompt examples that we think may be effective, if such prior knowledge is available. The goal is for the prompts to invoke various types of reasoning pathways when addressing queries, hence injecting diversity into the ensemble. Additional details are in Section A.1.
@@
    Prompt Selector. Then, an optimization process is performed over the candidate pool of prompts $\mathcal{W}$ to select a subset of $n$ prompts (i.e., $\{w_{i}\in\mathcal{W}\}_{i=1}^{n}$), based on a diversity metric that acts as an approximation of the relative performance of each subset (Section 3.1).
@@
    Response Aggregator. Finally, the responses from the $n$ constituent LLMs are aggregated through a response aggregator operation $\mathcal{A}$ to produce a single final response for the ensemble (Section 3.2).
@@
Putting everything together, our Dipper framework characterizes an ensemble of size $n$ via $\mathcal{E}(q_{t};M,n,\{w_{i}\}_{i=1}^{n})\coloneq\mathcal{A}(\{M(q_{t},w_{i})\}_{i=1}^{n})\rightarrow\hat{c}_{t}$, where the subset of prompts $\{w_{i}\}_{i=1}^{n}$ is chosen from a candidate pool $\mathcal{W}$ to optimize the expected ensemble performance $F(\mathcal{E},\mathcal{T})$ for a task $\mathcal{T}$.
3.1 Prompt Selector
With our framework, the optimization problem in Eq. 2 reduces to an optimization to choose the best subset of prompts $\{w_{i}\}_{i=1}^{n}$ from the set of candidate prompts $\mathcal{W}$:
$\operatorname{argmax}_{\{w_{i}\in\mathcal{W}\}_{i=1}^{n}}F(\mathcal{E}(q_{t};M,n,\{w_{i}\}_{i=1}^{n}),\mathcal{T}).$ (3)
Unfortunately, directly optimizing Eq. 2 is a combinatorial problem that is very challenging, even if a development/validation set is available for the task of interest. For example, selecting 5 prompts from a candidate pool of 200 prompts involves searching over (${200\choose 5}\approx 2.5\times 10^{9}$ candidates. Instead, we note that the best ensemble composition requires a balance of the two desiderata: fidelity and diversity. Hence, we propose optimizing Eq. 2 by considering how to prioritize the prompts that have the best predicted performance on the task $\mathcal{T}$, while maximizing the diversity of the selected set of prompts.
Prompt fidelity. First, we can approximate the predicted performance of each prompt by its average performance on a task development set $\mathcal{T}_{d}$ [^1] Without such a development set, an uninformed prior on the performance (e.g. uniform distribution across roles), or an informed-prior based on domain knowledge, could also be used.. Note that as inference using these various prompts on a small development set can be done in parallel, this process can in practice be significantly sped up by existing batch inference techniques such as those employed by vLLM [8]. Specifically, for a candidate pool of prompts $\mathcal{W}$ and development set $\mathcal{T}_{d}$, we can define a prompt fidelity mapping $u:\mathcal{W}\rightarrow[0,1]$,
$u(w)\coloneq F(M(\cdot,w),\mathcal{T}_{d}),$ (4)
where $M(\cdot,w)$ is the LLM model conditioned by prompt $w\in\mathcal{W}$, and $F$ the expected accuracy defined in Section 2. In practice, for a candidate pool of size $n$, $u(w)$ can be represented as an $n\times 1$ column vector, with the elements representing each prompt’s expected accuracy.
Semantic entropy. Instead, our approach involves prioritizing the prompts that have the best predicted performance on the task $\mathcal{T}$, while maximizing the diversity of the selected set of prompts. Then, we measure prompt diversity by considering how different the semantic meanings of the $n$ role prompts are from each other. We represent each prompt’s semantic meaning with a mapping $R$ from its text representation $w$ into a normalized continuous vector $s\in\mathbb{R}^{p}$ in a $p$-dimensional semantic embedding space $\mathcal{S}$ through a sentence embedding model $M_{s}$ [21], i.e., $R(w)\coloneq M_{s}(w)$. This mapping can be represented as an $n\times p$ prompt embedding matrix $R=[s_{1},\cdots,s_{n}]$ where $s$ is a $1\times p$ row vector representing each prompt.
To quantify prompt diversity of a given set of prompts, we propose to compute the volume enclosed by the selected prompts in semantic space. Intuitively, for $n$ fixed prompts, more diverse prompts point to more varied directions in semantic space, and enclose larger volume. Specifically, we define the semantic volume metric $V$ as
$V\coloneq\log\det(RR^{T}),$ (5)
where we take the logarithm (for numerical stability) of the Gram matrix determinant [^2] We omit a factor of 2 which does not affect the optimization process. For our setting, we also have $n<p$ as the semantic embedding space is usually high dimensional..
Fidelity-adjusted semantic volume. To incorporate the prompts’ expected accuracy information, we can compute the performance-adjusted prompt embedding matrix,
$\tilde{R}\coloneq\exp({\frac{\alpha}{2}\operatorname*{diag}(u)})R,$ (6)
where $\operatorname*{diag}(u)$ is the diagonal matrix with its $i^{\text{th}}$ diagonal element being the corresponding element $u_{i}$. This essentially scales each row $s_{i}$ in $R$ by an exponential factor based on its corresponding predicted accuracy, $\exp({\frac{\alpha}{2}u_{i}})$, where $\alpha$ is a scalar hyperparameter influencing the balance between diversity and expected performance. Intuitively, prompts with higher expected accuracy would then be able to support larger semantic volume and hence be prioritized for inclusion into the ensemble. The adjusted embedding matrix can then be used to compute the semantic volume in Eq. 5.
Optimization of semantic entropy. We can now recast Eq. 2 as an optimization of the fidelity-adjusted semantic volume metric $\tilde{V}$ evaluated over the set of candidate prompts. Note that instead of the expected ensemble performance $F(\mathcal{E})$, which is an objective that can only be optimized by blackbox optimization methods like Bayesian Optimization [22], our metric $V$ can be approximated by efficient, well-established heuristics.
Specifically, as the semantic volume metric is submodular, we can optimize for the best subset of roles by incrementally building the subset with a greedy approach up to the desired size $n$ and still be guaranteed a good approximation [23]. This allows us an efficient and theoretically-inspired approach to obtain the best ensemble prompts. Our full algorithm is outlined in Algorithm 1 in the Appendix.
3.2 Response Aggregator
Given the various constituent LLMs’ responses, the aggregation method determines how much information is used to derive the final ensemble output. We consider two approaches:
Majority voting (MV). The first involves extracting the final answer $\hat{c}$ from each LLM response $\hat{y}=\{\hat{r},\hat{c}\}$, and then selecting the answer that has been proposed the most number of times. This approach does not take into account the reasoning $\hat{r}$ output produced by the ensemble constituents, but is easily implementable.
LLM aggregation (LLMA). The second involves using another LLM instance to evaluate each constituent response, aggregate them, and generate a final answer for the task. This approach incurs additional LLM query cost and is dependent on the capabilities of the aggregator LLM, but has the advantage of potentially taking into account the various reasoning output $\hat{r}$ from the ensemble constituents to further improve overall performance (see Section 4.3 for details).
4 Experiments
Experimental set-up. We empirically evaluate our framework on mathematically reasoning tasks with the MATH [24], GSM8K, and MMLU-STEM datasets. We implement our framework by using the GPT-4o as our prompt generator and Qwen2-MATH-1.5B as the constituent model in the ensemble, where the ensemble constituents are run in parallel using vLLM [8] for fast batch inference. Further details of our experiments are in Appx. B.
Baselines. We evaluate our Dipper framework by comparing it against the "Self-ensemble" baseline, which lacks prompt diversity but incorporates diversity through repeated response sampling [15]. We also compare our Dipper implementation based on semantic volume ("Dipper") with two other variants: (1) a naïve implementation where prompts are sampled from the candidate pool based on their validation accuracy distribution ("Random+"), and (2) an ensemble using the "Top-n" prompts as evaluated on the validation set, which benefits from the diversity of prompts introduced by our prompt generation process but do not explicitly optimize for ensemble diversity otherwise.
4.1 Ensembles with fixed prompt methods
Figure 2: Plot of accuracy vs. average unique answers with ensembles of different numbers of prompts.
Figure 3: Comparison of different ensemble methods on MATH.
First, we illustrate the effectiveness of prompt-based ensemble methods by considering a fixed list of 7 reasoning prompts inspired by existing works [25, 26, 27] on prompting methods to boost reasoning capabilities (details in Appx. B.1). Under a fixed ensemble size of 7, Fig. 3 shows that the ensemble using the 7 different prompts (57.31%) significantly outperforms the self-ensemble with no prompt ((55.76%)) and the average performance (56.55%) of self-ensemble using any single prompt.
To further investigate the impact of prompt diversity, we evaluated all combinations of the 7 prompts while maintaining a fixed ensemble size of 7. For combinations with fewer than 7 prompts, we randomly sampled responses to reach a total of 7 before applying majority voting. The results in Fig. 3 reveals that increasing the number of prompts in the ensemble generally leads to higher accuracy, reduced variance, and fewer unique answers. The 7-prompt ensemble has the highest accuracy and lowest variance, which suggests that employing a diverse set of prompts in an ensemble can enhance performance and consistency, especially when we do not know which prompt would perform best before evaluation.
4.2 Ensembles with optimized prompt diversity
Next, we consider our full Dipper framework. We first generate a pool of prompt candidates ($|\mathcal{W}|=200$) using the 7 reasoning prompts in the previous section as in-context exemplars (details in Appx. B.1) and then perform diversity optimization (Sec. 3.1) to select the best ensemble prompts. Evaluation details are in Appx. B.2. As shown in Fig. 3, our method achieves the highest accuracy compared to all baseline ensemble methods across various ensemble sizes. Dipper also significantly outperforms the single LLM. For example, Dipper with $n=9$ has close to a 10%-pt increase (~20% accuracy gain) compared to the single LLM baseline. In fact, our ensemble that consists of just 3 Qwen2-MATH-1.5B model already slightly outperform the next model size class, the Qwen2-MATH-7B model. See more results on GSM8K and MMLU-STEM in Appx. C.2 where Dipper is shown to be consistently effective.
4.3 LLM aggregation can do better
Finally, we analyze the effects of using Majority voting (MV) or LLM aggregation (LLMA) for our response aggregator component (see experimental details in Appx. B.3). We consider ensembles of size $n=5$ with randomly selected prompts, and compare their performance on MATH when using either majority voting or LLM aggregation. Table 1 summarizes the results, showing that LLMA is more accurate than MV on average (i.e., higher $F(\mathcal{E})$). To better analyze the performance difference, we computed the “Override Ratio” which is how often a specific method is correct when the two methods disagree. Note that when MV and LLMA disagree, LLMA has a much higher ratio than MV which is only correct 8% of the time. We attribute LLMA’s advantage to its capability of understanding the reasoning $\hat{r}$ in responses even when the ensembles do not have a majority for the final answers $\hat{c}$. This is corroborated when we look at the number of unique answers $|C|$ when only one specific method is accurate: $|C|$ for LLMA is higher than that of MV, which suggests that LLMA performs better than MV when the ensemble produces more unique answers, as expected.
Method $F(\mathcal{E})$ Override Ratio $|C|$
MV 56.63 0.08 3.16
LLMA 64.87 0.29 3.70
Table 1: Comparison between MV and LLMA. $F(\mathcal{E})$ is the test performance. Override ratio is how often a specific method is correct when the two methods disagree. $|C|$ is the number of unique answers when only one specific method is accurate.
Figure 4: Comparison of Dipper and Dipper used together with self-reflection on MATH.
4.4 Dipper combined with other prompting methods like Reflexion
In addition, we also show that our ensemble framework Dipper is orthogonal to other established prompting techniques (e.g. CoT and Reflexion [6]), allowing it to stack and bring greater performance. In our experiments, we first use Dipper to select $n$ agents and query each agent with the questions. Their initial responses will be self-reflected according to the method proposed in Reflexion [6], before being aggregated into the final answer with MV. The results in Fig. 4 shows that Dipper coupled with reflection achieves much better results, suggesting that Dipper has the potential to be extended further or combined with other methods.
5 Conclusion
In this work, we have proposed a novel framework, Dipper, where a single LLM model type is fed an optimized, diverse set of reasoning prompts in parallel, effectively producing an ensemble at inference time to achieve performance improvement in reasoning tasks. Our empirical findings have demonstrated Dipper’s effectiveness in improving inference performance for a variety of reasoning tasks, which may inspire future works to investigate additional optimization methods for prompt-based inference-time ensembles to further improve performance gains.
Acknowledgments
This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD/2023-01-039J). This research is part of the programme DesCartes and is supported by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. This research is supported by the National Research Foundation Singapore and the Singapore Ministry of Digital Development and Innovation, National AI Group under the AI Visiting Professorship Programme (award number AIVP-$2024$-$001$).
Appendix A Additional details on the Dipper framework
A.1 Prompt Generator

### v2

@@ Abstract
Large Language Models (LLMs), particularly smaller variants, still struggle with complex reasoning tasks. While inference-time prompting can guide reasoning, existing methods often rely on sequential queries. Ensemble approaches offer a promising path to performance gains, especially given recent batch inference speed-ups. This work introduces Dipper, a novel, training-free framework that transforms a single LLM into an effective inference-time ensemble. By feeding the model an optimized and diverse set of prompts in parallel, Dipper elicits varied reasoning paths, leading to performance gains. We empirically demonstrate significant improvements on reasoning benchmarks, such as MATH, where a Dipper ensemble of three Qwen2-MATH-1.5B instances (via parallel prompting of a single model) outperforms a larger 7B model.
Dipper: Diversity in Prompts for Producing Large Language Model Ensembles in Reasoning Tasks
Wenyang Hu∗1,2, Gregory Kang Ruey Lauthanks: Equal contribution. 1,3, Diwen Liu1, Jizhuo Chen1, See-Kiong Ng1, Bryan Kian Hsiang Low1 1National University of Singapore, 2SAP, 3CNRS@CREATE, 1 Create Way, #08-01 Create Tower, Singapore 138602 {wenyang,greglau,lowkh}@comp.nus.edu.sg, seekiong@nus.edu.sg
@@ 1 Introduction
Despite remarkable advancements, Large Language Models (LLMs), particularly smaller models that are often constrained by resource limitations (e.g., GPU memory), continue to struggle with complex reasoning tasks (Huang and Chang, 2023). While inference-time methods offers a promising approach for enhancing LLM performance, especially for these smaller models (Snell et al., 2024), existing methods like Chain-of-Thought (CoT) and Reflexion often rely on sequential LLM queries, thereby incurring additional latency costs (Qiao et al., 2023; Zheng et al., 2023; Yao et al., 2023).
@@
Ensemble methods, which involve the use of multiple constituent models in parallel, have been shown to improve models’ performance and robustness in classical machine-learning settings (Ganaie et al., 2022) and are promising approaches to achieve better inference-time performance, although less well-studied in the LLM setting. The prospects of applying such methods to LLMs are increasingly attractive, given recent developments that have enabled significant speed-ups in parallel, LLM batch inference. These include methods to efficiently handle key-value cache memory (Kwon et al., 2023) and prompt caching to efficiently reuse common prompts for multiple queries (Zhu et al., 2024; Gim et al., 2024), enabling sub-linear (in the number of queries) costs for batch inference. However, a key challenge for successful ensembles is the diversity among their constituents (Krogh and Vedelsby, 1994; Zaidi et al., 2020). This principle extends to LLM ensembles, where achieving meaningful diversity from a single base model remains a central challenge.
Current approaches injecting such diversity, such as using heterogeneous model types (i.e., different LLMs) (Jiang et al., 2023; Huang et al., 2024), are often impractical due to memory constraints or use preferences for a single model type. Alternatively, methods like self-consistency, which rely on stochastic sampling from the same prompt (Wang et al., 2023b), typically yield limited diversity, thereby capping potential performance gains. We identify an influential yet overlooked source of diversity: system prompts. LLMs can generate varied reasoning pathways and outputs for the same task when guided by different instructional prompts (Kojima et al., 2023). This observation motivates our central research question: How can we systematically leverage prompt diversity to construct high-performing LLM ensembles from a single base model efficiently, without model retraining?
To address this, we introduce Dipper, a novel, training-free LLM ensemble framework that constructs an LLM ensemble by feeding a single base LLM an optimized, diverse set of reasoning prompts in parallel. This approach harnesses the parallel processing capabilities of modern LLM inference systems to achieve significant performance improvements in reasoning tasks, particularly for resource-constrained models. Dipper is notably simple, resource-efficient, and readily applicable to any black-box LLM via API access. Our key contributions are summarized as follows:
  • We propose Dipper, a novel framework for constructing inference-time ensembles from a single LLM using diverse reasoning prompts, adaptable to any (including black-box) LLM, and detail its core design principles (Sec. 4).
  • We develop a training-free, theory-inspired prompt diversity measure that, when used with our framework, can be efficiently optimized to maximize ensemble performance (Sec. 4.3).
  • We empirically demonstrate that our framework produces significant performance gains on math reasoning tasks (MATH, GSM8K, and MMLU-STEM), where our ensemble consisting of just a few small models (e.g., three Qwen2-MATH-1.5B) can outperform a larger model (e.g., Qwen2-MATH-7B) (Sec. 5).
2 Background and related works
LLMs and prompts.
@@ 2 Background and related works
Consider an LLM $M$ which can be viewed as a black box that encodes conditional probability distribution of text responses $y$ over any text input $q$ and prompt $w$, from which we can sample response $\hat{y}$, i.e.
@@
In practice, $w$ can be reasoning prompts that instruct LLMs to reason about $q$, e.g., "Let’s think step by step" in CoT (Wei et al., 2023). These prompts aim to influence the final LLM response, potentially by inducing additional LLM output (e.g., reasoning steps), and have been shown to yield performance boosts.
Prompt optimization.
To alleviate manual effort of prompt engineering, prompt optimization Zhou et al. (2023); Lin et al. (2024); Yang et al. (2024); Hu et al. (2024) works aim to automatically search for optimal prompts to maximize an LLM’s performance on specific tasks. However, such works have mainly focused on finding the best prompt for a single LLM, leaving the potential of optimizing prompts for LLM ensembles underexplored. In contrast, our work proposes a broader, novel framework for designing inference-time LLM ensembles with diverse prompts, and can incorporate existing prompt optimization methods. For example, we showed that an optimized prompt (e.g., "self-reflection" (Shinn et al., 2024)) can be combined with our framework.
LLM ensembles.
Ensemble methods, which combine multiple models to achieve superior performance and robustness (Ganaie et al., 2022), have seen limited application in the LLM domain. Prior LLM ensemble works have focused on heterogeneous ensembles that combine outputs from different LLM architectures or API providers (Jiang et al., 2023), multi-agent LLM settings that focus on interactions among agents (Du et al., 2023; Liu et al., 2023; Chen et al., 2023), or homogeneous self-ensembles that generate multiple responses from a single LLM using stochastic sampling (Wang et al., 2023b).
However, to the best of our knowledge, we are not aware of any work that has proposed forming and optimizing homogeneous LLM ensembles where their diversity is injected through varying reasoning prompts to constituents with the same underlying LLM model. Our work’s focus on such an approach exploits LLMs’ unique capabilities of generating diverse output given only changes to their prompts, allowing for a simple but effective method to achieve significant training-free boosts to LLM reasoning performance using inference-time compute, especially given recent developments in LLM batch inference methods (Kwon et al., 2023; Zhu et al., 2024; Gim et al., 2024).
3 Problem formulation
[…]
@@ 3 Problem formulation
Our objective is to design an ensemble framework with an appropriate design parameter $\phi$ such that given fixed $M$, $n$ and a small labeled development set, we can efficiently maximize Eq.˜2 by optimizing for $\phi$ to produce the best performing ensemble without additional training.
@@ 4 Method
A key driver of the ensemble’s performance is the diversity present among the constituents in the ensemble. Intuitively, a group where every member thinks the same way as each other is likely to result in less robust reasoning and decision-making (e.g. “groupthink”) compared to a group consisting of members with diverse thinking styles. Similarly, having an ensemble of LLM instances where all constituents are identical may be expected to yield less performance advantage compared to a diversified ensemble. In our setting where only a single LLM model is available, self-ensembles (Wang et al., 2023b) are examples of the former, as the constituents rely on LLM sampling stochasticity to generate potentially diverse responses, but will nonetheless still be sampling from the same distribution in Eq.˜1 and hence face limited diversity. Our framework for LLM ensembles, Dipper, efficiently introduces such diversity at inference time even when only one LLM model is available, through the use of high fidelity, diverse prompts.
In this section, we first provide an overview of our framework Dipper, before elaborating on the various components.
4.1 Overview of the Dipper framework
Drawing inspiration from how using different prompts $w$ would result in varying response distributions in Eq.˜1 given the same model $M$, our Dipper framework has the set of prompts $\{w_{i}\}_{i=1}^{n}$ fed into the ensemble of $n$ LLM instances as the key ensemble design parameter $\phi$.
Dipper consists of the following three components:
@@ 4.1 Overview of the Dipper framework
    Prompt Generator. First, an LLM generates a large candidate pool of prompts (denoted as $\mathcal{W}$), which can be based on some description of the task and in-context prompt examples that we think may be effective, if such prior knowledge is available. The goal is for the prompts to invoke various types of reasoning pathways when addressing queries, hence injecting diversity into the ensemble.
@@
    Prompt Selector. Drawing parallel to data/prompt selection Wu et al. (2024); Wang et al. (2025); Chen et al. (2025); Lau et al. (2024); Hemachandra et al. (2025), we select a subset of $n$ prompts $\{w_{i}\in\mathcal{W}\}_{i=1}^{n}$ from the candidate pool of prompts $\mathcal{W}$, where the selection is optimized based on a diversity metric that acts as an approximation of the relative performance of each subset.
@@
    Response Aggregator. Finally, the responses from the $n$ constituent LLMs are aggregated through a response aggregator operation $\mathcal{A}$ to produce a single final response for the ensemble.
@@
Putting everything together, our Dipper framework characterizes an ensemble of size $n$ via $\mathcal{E}(q_{t};M,n,\{w_{i}\}_{i=1}^{n})\coloneq\mathcal{A}(\{M(q_{t},w_{i})\}_{i=1}^{n})\rightarrow\hat{c}_{t}$, where the subset of prompts $\{w_{i}\}_{i=1}^{n}$ is chosen from a candidate pool $\mathcal{W}$ to optimize the expected ensemble performance $F(\mathcal{E},\mathcal{T})$ for a target task $\mathcal{T}$. We now describe each component in detail.
4.2 Prompt Generator